OpenStudy (miss.rose):

Linear Functions Part 2:

OpenStudy (kevin):

Is this about your last question?

OpenStudy (miss.rose):

Yes.

OpenStudy (miss.rose):

Yes, sorry for the confusion

OpenStudy (kevin):

No problem :) If you want to answer point b, this is what should you do: first, find the gradient using: m = (y2-y1)/ (x2-x1) second, find the linear equation using y2-y1 = m(x2-x1) third, check every variable if it's satisfied the linear equation, using y = weight x = dosage If you have do this all steps, you know the answer

OpenStudy (miss.rose):

Do you mean if you have to do all these steps, you know the answer?

OpenStudy (kevin):

yes, that's what I mean

OpenStudy (miss.rose):

okay, so the slope is 4.8. y=2.5 x=12 so it would be set up like this: 2.5=4.8(12)+b

OpenStudy (miss.rose):

Is that correct?

OpenStudy (kevin):

what is b?

OpenStudy (miss.rose):

b is the y-intercept. But I don't know what that is.

OpenStudy (kevin):

Hmm,, we supposed to answer point b not point a right?

OpenStudy (kevin):

or you want we to answer point a too

OpenStudy (miss.rose):

wait do you mean question a & b. Cause I just need question b answered.

OpenStudy (kevin):

alright, did you already answer question a?

OpenStudy (miss.rose):

yes

OpenStudy (kevin):

They have a same slope?

OpenStudy (miss.rose):

yes

OpenStudy (miss.rose):

Wait. the teacher said I don't need to answer that question. Sorry, Can you help with some other questions?

OpenStudy (kevin):

That would be the answer for the first question on point b: they relationship is linear because when we find the rate of change from one group to the next will get the same rate of change. So relation between weight and dosage gives a straight line when plotted on a graph.

OpenStudy (kevin):

If you want to answer the second question on point b, use formula: y-y1 = m(x-x1) with m is the slope, that is 4.8